OpenStudy (anonymous):

Evaluate the expression for x = - 6.1. 38.1 + x

OpenStudy (eric_d):

\(Welcome~to~Open~Study!!!\)

OpenStudy (eric_d):

substitute x into the expression

OpenStudy (anonymous):

substitute your value for x into the polynomial: 38.1 + x = 38.1 - 6.1 = You finish it up.

OpenStudy (anonymous):

@Steezysteve